A hiker walks up and down a hill. The hill has a cross section that can be modeled by y = -4/3|x - 600| + 800 where x and y are measured in feet and 0 ≤ x ≤ 1200. How far does the hiker walk?
1
Expert's answer
2020-09-17T14:07:18-0400

Answer:

The hiker walked a total distance of "2,000 \space meters"


Solutions


The cross section of the hill is given as

"y = -\dfrac {4}{3} \big|x - 600 \big| + 800" where "y" is the vertical distance and "x" is the horizontal displacement, and "0 \leq x \leq 120."The given equation for the cross section is a modulus function. Let us assume that "x \space and \space y" are measured in "meters (m)."


When "x = 0 \space m" , "y = -\dfrac {4}{3} \big|0-600 \big| + 800"

"= -\dfrac {4}{3} \big|-600 \big| + 800"

"= -\dfrac {4}{3}(600) + 800"


"= -800 + 800"


"= 0 m"

Thus, when "x = 0 m" the hiker was at the base of the hill.



Also, When "x = 120 m" ,

"y = -\dfrac {4}{3} \big| 120 - 800 \big| + 800"


"= -\dfrac {4}{3}(600) + 800"


"= -800 + 800"


"= 0 m"

Thus, when "x = 120 m" the hiker had descended from the hill top back to the base.


The maximum value of "y" occurs when "\big| x - 600 \big| = 0" , that is when "x = 600"

"\therefore \space max(y) = 800 m"


Therefore, the hiker ascended the hill with a horizontal displacement of 600m and a vertical displacement of 800m, and descended the hill with the same values.


Distance travelled

= distance up the hill + distance down the hill.

"= 2 × \sqrt {x^2 + y^2}"


"= 2× \sqrt {(600m)^2 + (800m)^2}"


"= 2 × \sqrt {360,000m^2 + 640,000m^2}"


"= 2 × \sqrt {1,000,000m^2}"


"= 2 \u00d7 1,000m"


"= 2,000 \space meters"


"\therefore" the hiker waked a distance of 1,000m up the slope and a distance of 1,000m down the slope giving a total distance walked of 2,000 meters.